OpenStudy (anonymous):

Direct/indirect variation: biologists count the total # of animals in a small section of a habitat. The total population is directly proportional to the size of the habitat. Write an equation to solve for the constant of variation 'k.'

OpenStudy (mathteacher1729):

This is translating an English sentence into a mathematical equation. All we're really doing is condensing notation (using fewer symbols). The total population is directly proportional to the size of the habitat. The total population = directly proportional to the size of the habitat. (is means "equals") Population(size of habitat) = directly proportional to the size of the habitat. ( Population is A FUNCTION OF the size of the habitat... meaning you input habitat size, your output is population). P(h) = directly proportional to the size of the habitat. (Abbreviating "size of habitat" to "h") P(h) = directly proportional to h (more abbreviation just like before) P(h) = k * h ("directly proportional to" means "some constant k times...(whatever it's directly proportional to) So the final equation is P(h) = k*h Or something like that. :)

OpenStudy (anonymous):

That's similar to what I had, but I was simply using different variables. I still don't quite understand direct and indirect variation.

OpenStudy (mathteacher1729):

Do you mean direct and INVERSE variation?

OpenStudy (anonymous):

Oh, yes. Sorry about that. There's a follow up question asking " A biologist counted 12 deer in a 100 acre parcel of land. Find the constant of variation 'k.'" I don't need the answer so much as I need to know how to work this.

OpenStudy (anonymous):

Wait, is the constant of variation going to be k = 3/25?

OpenStudy (mathteacher1729):

Oh, that has nothing to do with inverse variation though. P = 12 deer h = 100 acres 12 = k*100 solve for k

OpenStudy (anonymous):

Oh good! I did it right. Thanks so much!

OpenStudy (mathteacher1729):

Yup k = 3/25 , you got it.